caira12:

Solve for x: (5 points) (-6/5)X-4 = -46 Select one: a. 23 b. 35 c. −23 d. −35

Mercury:

hm (-6/5)X-4 = -46 since we have 4 subtracted from both sides, can you try adding 4 to both sides ?

NonportritV2:

(-6/5)X = -46+4=-42

NonportritV2:

(-6/5)X= -42

Mercury:

@caira12 still there? any ideas where to go from there?

caira12:

im a bit lost

Mercury:

no problem, let's try and break it down a bit ^^ (-6/5)X-4 = -46 since we have 4 subtracted from both sides, we simply add 4 to both sides to get (-6/5)X-4 + 4 = -46 + 4 make sense so far?

caira12:

yes

Mercury:

after we simplify that we get (-6/5)x = -42 (since -4 + 4 = 0 and -46 + 4 = -42)

Mercury:

now, since we have x multiplied by (-6/5) we can divide both sides by (-6/5) to get (-6/5)x / (-6/5) = -42/(-6/5) = what does -42 divided by (-6/5) equal?

caira12:

35

Mercury:

good ^^ so 35 = your sol'n

caira12:

thank you